Kabbed:

Solve the following system of equations and show all work. y = −x2 + 4 y = 2x + 1

Kabbed:

@dude help me please

dude:

Set them equal to each other

Kabbed:

okay

Kabbed:

-x^2 + 4 = 2x + 1 is this good

dude:

Yep now you want to get the whole equation equal to 0, do you know how to do that? \(-x^2 + 4 = 2x + 1 \)

Kabbed:

Move the equation to the left

dude:

Yes, can you do it?

Kabbed:

x^2 + 2x - 3 = 0

dude:

Good, now factor it

Kabbed:

I need help I don't know how to factor

dude:

Okay, what are the factors of 3?

Kabbed:

6

dude:

Factors D: Do you know what factors are?

Kabbed:

no

dude:

Ah D: You must know multiplication in order to do this Factors are the numbers that multiply to something The factors of 6 are: 1.6 2,3 1 times 6 is 6 2 times 3 is 6 Does that make sense?

Kabbed:

yes so 1 times 3 is 3 and nothing else so the factors are 1 and 3

dude:

Exactly now you want them to add up to 2 So which would you turn negative?

Kabbed:

3

dude:

Is -3+1 = 2?

Kabbed:

no

dude:

No, you turn the 1 negative 3-1 =2

dude:

Your terms would be \((x+3)(x-1)=0\) Do you know what to do from here?

Kabbed:

sorry no

Kabbed:

factor

Kabbed:

Factor more

dude:

Set each term equal to 0 (Split them) Important note: If you have factored it is very rare that you will need to factor again (Like 99% sure you wont need to unless you didn't factor completely) \((x+3)=0~~~~~and~~~~~~~~(x-1)=0\)

dude:

Now can you solve for x?

Kabbed:

sorry no

Kabbed:

dude where did you go

dude:

Boye (ಥ ͜ʖಥ) This is simple subtraction/addition.

dude:

What is the opposite of addition? What is the opposite of subtraction?

Kabbed:

yes I know the opposites of subtraction and addition

Kabbed:

but I still don't know what to do

563blackghost:

Do you know your negative and positive numbers?

Kabbed:

I don't know

563blackghost:

For example, Negative: -3,-2,-1 Positive: 3, 2, 1

Kabbed:

yes

Kabbed:

Now

Kabbed:

are you still here

563blackghost:

yes one second

Kabbed:

okay

563blackghost:

You have: \(\bf{x+3=0}\) What can you add to 3 to equal 0?

Kabbed:

-3

563blackghost:

Correct. So \(\bf{x=-3}\). \(\bf{x-1=0}\) What can you subtract 1 to equal 0?

Kabbed:

1

563blackghost:

So your solutions are: \(\bf{x=1,-3}\)

Kabbed:

I thought x had to be the same number for both of the problems thank you

563blackghost:

In some equations x can have multiple solutions.

563blackghost:

No problem